OpenStudy (day_2010):

How do you solve this: 1-3/2x=7/4

Parth (parthkohli):

=> -2/2x = 7/4 => -1/x = 7/4 => -1 = 7x/4 => -4 = 7x => x = -7/4

jhonyy9 (jhonyy9):

4x-6=7x -6=3x x=-6/3 x=-2

Parth (parthkohli):

OH, I misunderstood the question.

OpenStudy (day_2010):

Oh okay so where did you get the first part from jhonyy9? i think i understand what you did not too sure though?

jhonyy9 (jhonyy9):

so first of all you need to know that 4x will be the denominator common - yes ?

jhonyy9 (jhonyy9):

ok ?

OpenStudy (day_2010):

yes

OpenStudy (day_2010):

my computer froze?

jhonyy9 (jhonyy9):

now is understandably sure ?

OpenStudy (day_2010):

Yes sir,and thank you!

jhonyy9 (jhonyy9):

was my pleasure good luck bye